# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ Appendix B Output Layer Initialization
For categorical features $j$, we initialize the output layer such that the model achieves the respective losses under no information. Using the loss normalization constant $Z_{j}$ (see Appendix A) and dropping the expectation over $p(\epsilon)$, we have
@@
Appendix C Adaptive Normalization of the Average Diffusion Loss
Both the loss calibration (see Appendix A) and output layer initialization (see Appendix B) ensure that the losses across timesteps (and features) are equal at initialization. During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest. However, the better the model becomes at a given timestep $t$, the lower the loss at the respective timestep, and the lower the gradient signal relative to the signal for timesteps $\tilde{t}>t$. We counteract this with adaptive normalization of the average diffusion loss (averaged over the features) across timesteps. Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 6, such that the normalized loss is the same (equal to one) for all $t$. Similar methods have been used by Karras et al. (2023) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and use the MSE loss to learn this weighting. First, we compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022). Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features. The result is passed through a single linear layer to output a scalar value, passed through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$. We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.
Appendix D Derivation of the Functional Timewarping Form
@@
Appendix E Benchmark Datasets
Our selected benchmark datasets are highly diverse, particularly in the number of categories for categorical features (see Table 3). For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for ease of presentation. For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area. In the diabetes data, we convert the task by predicting whether a patient was readmitted to a hospital. All datasets are publicly accessible and (except nmes) licensed under creative commons.
Table 3: Overview of the selected experimental datasets. We count the outcome towards the respective features that remain after removing continuous features with an excessive number of missings. The minimum and maximum number of categories are taken over all categorical features.
@@ Appendix E Benchmark Datasets
observations categorical continuous    min. max.
@@
adult (Becker & Kohavi, 1996) CC BY 4.0 binary classification 48 842 9 6 2 42
bank (Moro et al., 2012) CC BY 4.0 binary classification 41 188 11 10 2 12
beijing (Chen, 2017) CC BY 4.0 regression 41 757 1 10 4 4
churn (Keramati et al., 2014) CC BY 4.0 binary classification. 3 150 5 9 2 5
covertype (Blackard, 1998) CC BY 4.0 binary classification 581 012 44 10 2 2
default (Yeh, 2016) CC BY 4.0 binary classification 30 000 10 14 2 11
diabetes (Clore et al., 2014) CC BY 4.0 binary classification 101 766 28 9 2 716
lending (Club, 2015) DbCL 1.0 regression 9 182 10 34 2 3151
@@
Appendix F Baseline Models
@@
(Zhang et al., 2024) – a diffusion-based model that first learns a tranformer-based VAE to map mixed-type data to a continuous latent space. Then, the diffusion model is trained on that latent space. We use the official code available at https://github.com/amazon-science/tabsyn under the Apache 2.0 license.
Appendix G Implementation Details
@@
TabSyn 3 001 646
CDTD (per type schedule, TabDDPM architecture) 2 999 721
We adjust each architecture to a total of ±$\pm 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets. Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the onehot encoding but is still comparable across models. We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM, TabSyn and CDTD to 256. To align TabDDPM, TabSyn and CDTD further, we use the TabDDPM architecture for all models, with appropriate adjustments for different input types and dimensions. If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
@@
(ours): To ensure comparability in particular to TabDDPM, CoDi and TabSyn, we use the same neural network architecture as TabDDPM. We only change the input layers to accommodate our embedding-based framework. In the input layer, we vectorize all embedded categorical features and concatenate them with the scalar valued continuous features. The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature. Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 808 neurons per layer we use 806) on the adult dataset. More details on the CDTD implementation are given in Appendix J.
Appendix H Tuning of the Detection Model
@@
Appendix I Machine Learning Efficiency Models
@@
Appendix J CDTD Implementation Details
@@
To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture and optimizer as Kotelnikov et al. (2023), which was also adapted by TabSyn (Zhang et al., 2024). An overview of the score model is provided in Figure 4: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, are projected onto a 256-dimensional space. Similarly, timestep $t$ and possibly conditioning information $y$ are embedded in the same space. Then, all 256-dimensional vectors are added and the results is processed by a set of five fully-connected linear layers with ReLU activation functions. Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
@@
The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space. The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont},i}$, and $\lvert C_{j}\rvert$ values for each $x_{\text{cat},j}$, with $C_{j}$ the set of categories of feature $j$. We change the initialization of the output layer as described in Appendix B: To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero. We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022). Also, Dieleman et al. (2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{init}}$, is an important hyperparameter. In this paper, we set $\sigma_{\text{init}}=0.001$ for all datasets and have not seen detrimental effects. Table 7 indicates that CDTD is not sensitive to the choice of $\sigma_{\text{init}}$.
@@
Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 808 for TabDDPM to 806). Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we scale our timesteps $t\in[0,1]$ up by 1000. We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000). However, since we work with embeddings we add a linear warmup schedule over the first 100 steps.
Instead of using the vanilla uniform (time)step sampling as the TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2022). The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules. For generation, we use an Euler sampler with 200 steps to minimize the discretization error.
Appendix K CDTD Sampling
@@
To sample from our learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2). For example, for two different features $x_{1}$ and $x_{2}$, we deconstruct the ODE as:
$d{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})dt$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}dt$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}dt$
@@
In practice, we use an Euler sampler with 200 discrete timesteps $\Delta t=t_{i+1}-t_{i}<0$. The timesteps are generated as a linearly spaced grid on $[0,1]$ and transformed afterwards into noise levels $\sigma_{k}(t)$ via the described timewarping procedure. For the discretized and simplified the ODE above, this yields
${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t)}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t)}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix}.$
@@
where $\odot$ denotes the element-wise product. Hence, we are effectively taking feature-specific steps of length $\Delta\sigma_{k}(t)$. The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{i}(t)$ is small).
@@
We use finite differences to approximate $\dot{\sigma}_{i}$, instead of the available, analytical variant, since $\frac{d\sigma_{k}(t)}{dt}\rightarrow\infty$ as $t\rightarrow 1$. The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx dt$ holds. For a large number of steps, this assumption does not hold in practice, and for $\frac{d\sigma_{k}(t)}{dt}$ the update of ${\mathbf{x}}$ overshoots the target drastically. Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{d\sigma_{k}(t)}{dt}$ into a good approximation of the actual change in $\sigma_{k}(t)$. Moreover, the analytical solution would approximate $d\sigma_{k}(t)=\dot{\sigma}_{k}(t)dt$, i.e., the change in the noise level caused by a change in $t$. Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $d\sigma_{k}(t)=\Delta\sigma_{k}(t)$. Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
@@
200 (default) 0.033 0.015 0.004 0.127 0.560 0.013 0.003 0.372
500 0.028 0.018 0.005 0.130 0.565 0.012 0.003 0.372
1000 0.028 0.018 0.005 0.129 0.560 0.012 0.003 0.373
1500 0.028 0.018 0.005 0.129 0.561 0.012 0.003 0.374
Appendix L Sensitivity to Important Hyperparameters
The training and sampling processes of CDTD are affected by various novel hyperparameters. Generally, a per-type noise schedule works best as we show in our main results in Table 1 for a diverse set of benchmark datasets. Here, we examine the sensitivity of CDTD to two additional important hyperparameters: (1) the standard deviation of the noise used to initialize the embeddings (and therefore specific to score interpolation), $\sigma_{\text{init}}$, and (2) the weight of the low noise levels used to initialize the $\mu_{k}$ in the adaptive noise schedule parameterization.
The experiments in Dieleman et al. (2022) show that $\sigma_{\text{init}}$ is a crucial hyperparameter for score interpolation on text data. The same sensitivity does not translate to the tabular data domain, as shown in our results in Table 7. The much smaller embedding dimension (16 vs. 256) and the feature-specific embeddings significantly decrease the number of distinguishable categories. Compared to a vocabulary size of 32000 for text data (Dieleman et al., 2022), we only face a maximum of 3151 categories in the lending dataset (see Table 3). Thus, unlike other generative (diffusion) models for tabular data, CDTD scales to a practically arbitrary number of categories.
Our proposed functional form for the adaptive noise schedules (see Appendix D) is the first to allow for the incorporation of prior information about the importance of low vs. high (normalized) noise levels. For this, we adjust the weight of low noise levels which directly determines the location of the inflection point $\mu_{k}$ (see Section 3.3). The results in Table 8 indicate low sample quality sensitivity to weight changes for a per-type noise schedule. The initialization only impacts the time to convergence but not (much) the location of the optimum. In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge.
@@
1 0.036 0.015 0.004 0.143 0.651 0.015 0.003 0.313
2 0.030 0.014 0.005 0.147 0.651 0.014 0.003 0.352
3 (default) 0.033 0.015 0.004 0.154 0.651 0.013 0.004 0.366
4 0.034 0.019 0.005 0.148 0.656 0.013 0.004 0.370
Appendix M Advantages of Diffusion in Data Space

### v4

@@ Appendix C Output Layer Initialization
For categorical features $j$, we initialize the output layer such that the model achieves the respective losses under no information. Using the loss normalization constant $Z_{j}$ (see Appendix B) and dropping the expectation over $p(\epsilon)$, we have
@@
Appendix D Adaptive Normalization of the Average Diffusion Loss
Both the loss calibration (see Appendix B) and output layer initialization (see Appendix C) ensure that the losses across timesteps (and features) are equal at initialization. During training, the adaptive noise schedules allow the model to focus automatically on the noise levels that matter most, i.e., where the loss increase is steepest. However, the better the model becomes at a given timestep $t$, the lower the loss at the respective timestep, and the lower the gradient signal relative to the signal for timesteps $\tilde{t}>t$. We counteract this with adaptive normalization of the average diffusion loss (averaged over the features) across timesteps. Specifically, we want to weight the average diffusion loss at timestep $t$, $\mathcal{L}(t)$ given in Equation 5, such that the normalized loss is the same (equal to one) for all $t$. Similar methods have been used by Karras et al. (2024) and Kingma & Gao (2023), we follow the latter in the setup of the corresponding network.
We train a neural network alongside our diffusion model to predict $\mathcal{L}(t)$ based on $t$ and use the MSE loss to learn this weighting. First, we compute $c_{\text{noise}}(t)=\log(t)/4$ following the EDM parameterization (Karras et al., 2022). Then, we embed $c_{\text{noise}}$ in frequency space (1024-dimensional) using Fourier features. The result is passed through a single linear layer to output a scalar value and through an exponential function to ensure that the prediction ^$\hat{\mathcal{L}}(t)\geq 0$. We initialize the weights and biases to zero, to ensure that at model initialization we have a unit normalization.
Appendix E Derivation of the Functional Timewarping Form
@@
Appendix F Benchmark Datasets
Our selected benchmark datasets are highly diverse, particularly in the number of categories for categorical features (see Table 3). For the diabetes and covertype datasets, we transform the original multi-class classification problem into a binary classification task for ease of presentation. For the diabetes data, we convert the task to predicting whether a patient was readmitted to a hospital. For the covertype data, the task is converted into predicting whether a forest of type 2 is present in a given 30 $\times$ 30 meter area. All datasets are publicly accessible and (except nmes) licensed under creative commons.
Table 3: Overview of the selected experimental datasets. We count the target towards the respective features that remain after removing continuous features with an excessive number of missings. The minimum and maximum number of categories are taken over all categorical features.
@@ Appendix F Benchmark Datasets
observations categorical continuous    min max
@@
adult (Becker & Kohavi, 1996) CC BY 4.0 binary class. 48 842 9 6 2 42
bank (Moro et al., 2014) CC BY 4.0 binary class. 41 188 11 10 2 12
beijing (Chen, 2015) CC BY 4.0 regression 41 757 1 10 4 4
churn (Keramati et al., 2014) CC BY 4.0 binary class. 3 150 5 9 2 5
covertype (Blackard, 1998) CC BY 4.0 binary class. 581 012 44 10 2 2
default (Yeh, 2009) CC BY 4.0 binary class. 30 000 10 14 2 11
diabetes (Clore et al., 2014) CC BY 4.0 binary class. 101 766 28 9 2 716
lending (Lending Club, 2015) DbCL 1.0 regression 9 182 10 34 2 3151
@@
Appendix G Baseline Models
@@
(Zhang et al., 2024) – a diffusion-based model that first learns a transformer-based VAE to map mixed-type data to a continuous latent space. Then, the diffusion model is trained on that latent space. Note that despite TabSyn utilizing a separately trained encoder, this does not result in a lower-dimensional latent space and therefore, does not speed up sampling. We use the official code available at https://github.com/amazon-science/tabsyn under the Apache 2.0 license.
Appendix H Implementation Details
@@
TabSyn 2 997 765
CDTD (per type) 3 002 969
We adjust each architecture to a total of $\sim 3$ million trainable parameters on the adult dataset to improve the comparability further (see Table 4) and use the same architectures for all considered datasets. Note that the total number of parameters may vary slightly across datasets due to different number of features and categories affecting the one-hot encoding but is still comparable across models. We also align the embedding/bottleneck dimensions for CTGAN, TVAE, TabDDPM, TabSyn and CDTD to 256. To align TabDDPM, TabSyn and CDTD further, we use the TabDDPM architecture for all models, with appropriate adjustments for different input types and dimensions. If applicable, all models are trained for 30k steps on a single RTX 4090 instance, using PyTorch version 2.2.2.
@@
(ours): To ensure comparability in particular to TabDDPM, CoDi and TabSyn, we use the same neural network architecture as TabDDPM. We only change the input layers to accommodate our embedding-based framework. In the input layer, we vectorize all embedded categorical features and concatenate them with the scalar valued continuous features. The adjusted output layer ensures that we predict a single value for each continuous features and set of class-specific probabilities for each categorical feature. Since our use of embeddings introduces additional parameters, we scale the hidden layers slightly down relative to the TabDDPM to ensure approximately 3 million trainable parameters (instead of 798 neurons per layer we use 796) on the adult dataset. More details on the CDTD implementation are given in Appendix K.
Appendix I Tuning of the Detection Model
@@
Appendix J Machine Learning Efficiency Models
@@
We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load.
Appendix K CDTD Implementation Details
@@
To enable a fair comparison to the other methods, and to TabDDPM and TabSyn in particular, the CDTD score model utilizes the exact same architecture as Kotelnikov et al. (2023), which was also adapted by TabSyn (Zhang et al., 2024). We use the QuantileTransformer to pre-process continuous features, followed by a standardization to zero mean and unit variance. An overview of the score model is provided in Figure 5: First, the noisy data, i.e., the noisy scalars for continuous features and the noisy embeddings for categorical features, and the timestep $t$, are projected onto a 256-dimensional space. Then, all 256-dimensional vectors are added and the result is processed by a set of five fully-connected linear layers with ReLU activation functions. Lastly, a linear projection maps the output of the fully-connected layers to the required output dimensions, which depend on the number of features and number of categories per feature.
@@
The only major difference to the TabDDPM setup are the inputs, as we need to embed the categorical features in Euclidean space. The output dimensions are the same, as we need to predict a single scalar for each $x_{\text{cont}}^{(i)}$, and $C_{j}$ values for each $x_{\text{cat}}^{(j)}$, with $C_{j}$ being the number of categories of feature $j$. We change the initialization of the output layer as described in Appendix C. To handle our inputs, we embed the categorical features in 16-dimensional space and add a feature-specific bias of the same dimension, which captures feature-specific information common to all categories and is initialized to zero. We L2-normalize each embedding to prevent a degenerate embedding space in which embeddings are pushed further and further apart (see Dieleman et al., 2022). Also, Dieleman et al. (2022) argue that the standard deviation of the Normal distribution used to initialize the embeddings, denoted by $\sigma_{\text{init}}$, is an important hyperparameter. In this paper, we set $\sigma_{\text{init}}=0.001$ for all datasets and have not seen detrimental effects. Table 7 indicates that CDTD is not sensitive to the choice of $\sigma_{\text{init}}$.
@@
Since we utilize embeddings, we have to scale the neurons per layer slightly down in the stack of the five fully-connected layers (from 798 for TabDDPM to 796). Also, since TabDDPM samples discrete steps from $[0,T]$, with $T\gg 1$, we scale our timesteps $t\in[0,1]$ up by 1000. We use the same optimizer (Adam), learning rate (0.001), learning rate decay (linear), EMA decay (0.999), and training steps (30000). However, since we work with embeddings we add a linear warmup schedule over the first 1000 steps. Lastly, instead of using uniform (time)step sampling as TabDDPM, the CDTD model uses antithetic sampling (Dieleman et al., 2022; Kingma et al., 2021). The timesteps are still uniformly distributed but spread out more evenly over the domain, which benefits the training of the adaptive noise schedules. For generation, we use a deterministic (Euler) sampler with 200 steps to minimize the discretization error (see Appendix L for details).
Appendix L CDTD Sampling
Algorithm 1 Deterministic Sampling
0:  $N$ sampling steps Sample ${\mathbf{x}}_{\text{cont},t_{0}}^{(i)}\sim\mathcal{N}({\bm{0}},\sigma_{\text{cont},\max}^{2}{\mathbf{I}}_{K_{\text{cont}}})\,\forall i\in\{1,\dots,K_{\text{cont}}\}$ Sample ${\mathbf{x}}_{\text{cat},t_{0}}^{(j)}\sim\mathcal{N}({\bm{0}},\sigma_{\text{cat},\max}^{2}{\mathbf{I}}_{K_{\text{cat}}})\,\forall j\in\{1,\dots,K_{\text{cat}}\}$
1:  for $s\in\{0,\dots,N-1\}$ do
2:     $t_{s}=1-s/N$
3:     $t_{s+1}=1-(s+1)/N$
4:     ${\mathbf{x}}_{s}\leftarrow({\mathbf{x}}_{\text{cat},t_{s}}^{(1)},\dots,{\mathbf{x}}_{\text{cat},t_{s}}^{(K_{\text{cat}})},x_{\text{cont},t_{s}}^{(1)},\dots,x_{\text{cont},t_{s}}^{(K_{\text{cont}})})$
5:     for all $i\in\{1,\dots,K_{\text{cont}}\}$ do
6:        $\mathrm{d}x_{\text{cont}}^{(i)}=(x_{\text{cont},t_{s}}^{(i)}-D^{\text{cont},i}_{\bm{\theta}}({\mathbf{x}}_{s},t_{s}))/\sigma_{\text{cont},i}(t_{s})$
7:        $x_{\text{cont},t_{s+1}}^{(i)}\leftarrow x_{\text{cont},t_{s}}^{(i)}+[\sigma_{\text{cont},i}(t_{s+1})-\sigma_{\text{cont},i}(t_{s})]\,\mathrm{d}x_{\text{cont}}^{(i)}$
8:     end for
9:     for all $j\in\{1,\dots,K_{\text{cat}}\}$ do
10:        $\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{s}}^{(j)})=D^{\text{cat},j}_{\bm{\theta}}({\mathbf{x}}_{s},t_{s})$
11:        $\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\Bigl{(}{\mathbf{x}}_{\text{cat},t_{s}}^{(j)}-\mathbb{E}_{\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{s}}^{(j)})}[{\mathbf{x}}_{\text{cat},0}^{(j)}]\Bigr{)}/\sigma_{\text{cat},j}(t_{s})$ from Equation 4
12:        ${\mathbf{x}}_{\text{cat},t_{s+1}}^{(j)}\leftarrow{\mathbf{x}}_{\text{cat},t_{s}}^{(j)}+[\sigma_{\text{cat},j}(t_{s+1})-\sigma_{\text{cat},j}(t_{s})]\,\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}$
13:     end for
14:  end forRecover classes from embeddings with additional pass to score model
15:  ${\mathbf{x}}_{N}\leftarrow({\mathbf{x}}_{\text{cat},t_{N}}^{(1)},\dots,{\mathbf{x}}_{\text{cat},t_{N}}^{(K_{\text{cat}})},x_{\text{cont},t_{N}}^{(1)},\dots,x_{\text{cont},t_{N}}^{(K_{\text{cont}})})$
16:  for all $j\in\{1,\dots,K_{\text{cat}}\}$ do
17:     $\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}|{\mathbf{x}}_{\text{cat},t_{N}}^{(j)})=D^{\text{cat},j}_{\bm{\theta}}({\mathbf{x}}_{N},t_{N-1})$
18:     $x_{\text{cat}}^{(j)}\leftarrow\operatorname*{arg\,max}_{c\in C_{j}}\hat{\mathrm{P}}({\mathbf{x}}_{\text{cat},0}^{(j)}={\mathbf{e}}_{c}^{(j)}|{\mathbf{x}}_{\text{cat},t_{N}}^{(j)})$ (pick most likely category)
19:  end for
20:  return  $(x_{\text{cat}}^{(1)},\dots,x_{\text{cat}}^{(K_{\text{cat}})},x_{\text{cont},t_{N}}^{(1)},\dots,x_{\text{cont},t_{N}}^{(K_{\text{cont}})})$
Algorithm 1 shows our deterministic sampling approach, where $D^{\text{cont},i}_{\bm{\theta}}$ and $D^{\text{cat},j}_{\bm{\theta}}$ represent the score model output for the $i$-th continuous and $j$-th categorical feature, respectively. We found results similar to Karras et al. (2022), i.e., that adding stochasticity to the sampling process does not benefit sample quality. We also experimented with second-order samplers (e.g., Heun) but found them to add little benefit over our first-order method while requiring double the NFEs. As a default, we use 200 sampling steps. Table 6 shows that the gains in sample quality are marginal to non-existent after more than 500 sampling steps.
There is another subtlety in our sampler: Unlike in EDM (Karras et al., 2022), in the final step the sampler steps into $t=0$, which is not associated with $\sigma_{0}=0$ but $\sigma_{0}=\sigma_{\text{min}}$. This is a consequence of utilizing more than one noise schedule. We condition the score model not on $\sigma$ but on $t$, which indicates the “global” time common to all noise schedules. Moving from $\sigma_{0}=\sigma_{\text{min}}$ to $\sigma=0$ in the final step, would imply $t<0$. Therefore, we let $\sigma_{\text{min}}=0$ for all feature types.
@@
To sample from the learned distribution, we need to run the reverse process of the probability flow ODE (Equation 2). For example, for two different features $x_{1}$ and $x_{2}$, we deconstruct the ODE as:
$\mathrm{d}{\mathbf{x}}$ $=-\frac{1}{2}{\mathbf{G}}(t){\mathbf{G}}(t)^{\mathsf{T}}\nabla_{\mathbf{x}}\log p_{t}({\mathbf{x}})\mathrm{d}t$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)\sigma_{1}(t)&\\ &\dot{\sigma_{2}}(t)\sigma_{2}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)^{2}}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)^{2}}\end{bmatrix}\mathrm{d}t$
$=-\begin{bmatrix}\dot{\sigma_{1}}(t)&\\ &\dot{\sigma_{2}}(t)\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t)}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t)}\end{bmatrix}\mathrm{d}t$
@@
In practice, we use an Euler sampler with 200 discrete timesteps $\Delta t=t_{i+1}-t_{i}<0$. The timesteps are generated as a linearly spaced grid on $[0,1]$ and transformed afterwards into noise levels $\sigma_{k}(t)$ via the described timewarping procedure. For the discretized and simplified ODE at step $i$, this yields
${\mathbf{x}}_{i+1}$ $={\mathbf{x}}_{i}-\begin{bmatrix}\frac{\Delta\sigma_{1}(t)}{\Delta t}&\\ &\frac{\Delta\sigma_{2}(t)}{\Delta t}\end{bmatrix}\begin{bmatrix}\frac{\hat{x_{1}}-x_{1}}{\sigma_{1}(t_{i})}\\ \frac{\hat{x_{2}}-x_{2}}{\sigma_{2}(t_{i})}\end{bmatrix}\Delta t={\mathbf{x}}_{i}+\begin{bmatrix}\frac{x_{1}-\hat{x_{1}}}{\sigma_{1}(t_{i})}\\ \frac{x_{2}-\hat{x_{2}}}{\sigma_{2}(t_{i})}\end{bmatrix}\odot\begin{bmatrix}\Delta\sigma_{1}(t)\\ \Delta\sigma_{2}(t)\end{bmatrix},$
@@
where $\odot$ denotes the element-wise product. Hence, we are effectively taking feature-specific steps of length $\Delta\sigma_{k}(t)$. The adaptive noise schedules (timewarping) therefore not only affect the training process, but also focus most work in the reverse process on the noise levels that matter most for sample quality (i.e., where $\Delta\sigma_{k}(t)$ is small).
@@
We use finite differences to approximate $\dot{\sigma}_{i}$, instead of the available, analytical variant, since $\frac{d\sigma_{k}(t)}{dt}\rightarrow\infty$ as $t\rightarrow 1$. The step $\Delta t$ would therefore be required to decrease as $t\rightarrow 1$ to ensure $\Delta t\approx\mathrm{d}t$ holds. For a large number of steps, this assumption does not hold in practice, and for $\frac{\mathrm{d}\sigma_{k}(t)}{\mathrm{d}t}$ the update of ${\mathbf{x}}$ overshoots the target drastically. Intuitively, $\sigma_{k}(t)$ becomes too steep near the terminal timestep $t=1$ such that the step size can not sufficiently compensate for the slope increase to turn $\frac{\mathrm{d}\sigma_{k}(t)}{\mathrm{d}t}$ into a good approximation of the actual change in $\sigma_{k}(t)$. Moreover, the analytical solution would approximate $\mathrm{d}\sigma_{k}(t)=\dot{\sigma}_{k}(t)\mathrm{d}t$, i.e., the change in the noise level caused by a change in $t$. Since we know exactly where $\sigma_{k}(t)$ will end up when changing $t$, we are better off using that exact value and let $\mathrm{d}\sigma_{k}(t)=\Delta\sigma_{k}(t)$.
@@
100 0.038 0.011 0.007 0.131 0.574 0.012 0.003 0.315
200 (default) 0.030 0.009 0.006 0.130 0.565 0.012 0.003 0.316
500 0.027 0.009 0.006 0.130 0.557 0.013 0.002 0.313
1000 0.028 0.008 0.006 0.130 0.562 0.013 0.002 0.311
Appendix M Sensitivity to Important Hyperparameters
The training and sampling processes of CDTD are affected by various novel hyperparameters. Generally, a per-type noise schedule works best, as we show in our main results in Table 1 for a diverse set of benchmark datasets. Here, we examine the sensitivity of CDTD to two additional important hyperparameters: (1) the standard deviation of the noise used to initialize the embeddings (and therefore specific to score interpolation), $\sigma_{\text{init}}$, and (2) the weight of the low noise levels used to initialize the $\mu_{k}$ in the adaptive noise schedule parameterization.
The experiments in Dieleman et al. (2022) show that $\sigma_{\text{init}}$ is a crucial hyperparameter for score interpolation on text data. Table 7 shows that this sensitivity does not translate to the tabular data domain. This may be explained by the much smaller embedding dimension (16 vs. 256) or by our usage of feature-specific embeddings. Compared to a vocabulary size of 32000 for text data (Dieleman et al., 2022), we only face a maximum of 3151 categories in the lending dataset (see Table 3). Thus, unlike other generative (diffusion) models for tabular data, CDTD scales to a practically arbitrary number of categories.
Our proposed functional form for the adaptive noise schedules (see Appendix E) is the first to allow for the incorporation of prior information about the importance of low vs. high (normalized) noise levels. For this, we adjust the weight of low noise levels which directly determines the location of the inflection point $\mu_{k}$ (see Section 3.3). The results in Table 8 indicate low sensitivity of sample quality to weight changes for a per-type noise schedule. The initialization only impacts the time to convergence but not (much) the location of the optimum. In our experiments, the number of training steps (30000) appears to be high enough for all model variants to converge to similar states.
@@
1 0.031 0.010 0.005 0.126 0.570 0.012 0.002 0.238
2 0.030 0.009 0.004 0.125 0.570 0.013 0.002 0.289
3 (default) 0.030 0.009 0.006 0.130 0.565 0.012 0.003 0.316
4 0.028 0.011 0.005 0.134 0.574 0.011 0.003 0.350
Appendix N Advantages of Diffusion in Data Space